OpenStudy (anonymous):

suppose csc x=5, cos x > 0, find cos x

OpenStudy (anonymous):

you know \[\csc(x)=5\] so right away you know that \[\sin(x)=\frac{1}{5}\]

OpenStudy (anonymous):

now the easiest thing to do is draw a right triangle with "opposite" side 1 and hypotenuse 5. the "adjacent" side will be \[\sqrt{5^2-1^2}=\sqrt{24}\] so \[\cos(x)=\frac{\sqrt{24}}{5}\]

OpenStudy (anonymous):

you might want to write \[\sqrt{24}=2\sqrt{6}\]